# Re-biologizing RL: The relevance of Darwinian neurodynamics and origin-of-life to reinforcement-learning

### v1

@@ Abstract
Three core tenets of reinforcement learning (RL)—concerning the definition of agency, the objective of learning, and the scope of the reward hypothesis—have been highlighted as key targets for conceptual revision, with major implications for theory and application. We propose a framework, inspired by open-ended evolutionary theory, to reconsider these three "dogmas." We revisit each assumption and address related concerns raised alongside them. To make our arguments relevant to RL as a model of biological learning, we first establish that evolutionary dynamics can plausibly operate within living brains over an individual’s lifetime, and are not confined to cross-generational processes. We begin by revisiting the second dogma, drawing on evolutionary insights to enrich the “adaptation-rather-than-search” view of learning. We then address the third dogma regarding the limits of the reward hypothesis, using analogies from evolutionary fitness to illuminate the scalar reward vs. multi-objective debate. After discussing the implications for exploration in RL, we turn to the first—and arguably most fundamental—issue: the absence of a formal account of agency. We argue that unlike the other two problems, the evolutionary paradigm alone cannot resolve the agency question, though it gestures in a productive direction. We advocate integrating ideas from origins-of-life theory, where the thermodynamics of sustenance and replication offer promising foundations for understanding agency and resource-constrained reinforcement learning in biological systems.
@@ 1.1 Motivation & Overview
Finally, in section 3.2, we take a more practical turn and highlight areas where this theoretical framing can provide new tools for a practicing RL engineer. We specifically focus on the implications of the evolutionary framework on the problems of exploration-exploitation and resource-constrained reinforcement learning.
@@ 1.2 Evolution in the Brain
One of the leading figures in establishing the clonal selection theory of learning by the immune system, Gerald Edelman, was also one of the early advocates of a similar mechanism of learning by the nervous system (Edelman, 1993). This research program came to be known as neural Darwinism, and despite complementary ideas in a similar vein (Calvin, 1998), its adoption has been limited. While the volume of research on experience-guided learning via neuro-plasticity has catapulted over the years, research on the neural mechanisms and correlates of such neural Darwinism remain sparse.
@@ 2.1 Adaptation Versus Search
We will revisit the second point in section 2.2 when we discuss fitness in relation to scalar and multi-objective optimization. It is however constructive to elaborate on the first point and explain the mechanism with which novelty and diversity are achieved in these algorithms. Instead of “selection pressure”, the biological inspiration for novelty search is the notion of “speciation” or “niching” (Stanley & Miikkulainen, 2002; Mahfoud, 1996). This concept captures the tendency of organisms to found a new niche and thereby “avoid competition and exploit untapped resources” (Lehman & Stanley, 2010). QD algorithms encourage a diverse “archive” of solutions, by limiting selection pressure to individuals within the same niche, and avoiding premature selection against candidates that inhabit novel behavioral or phenotypic characteristics. An adjacent concept called “niche construction” (Odling-Smee et al., 2003; 2013) further enriches the adaptational view of learning in contrast to the search metaphor: organisms not only occupy a niche, but actively modify them. This results in coevolutionary dynamics between a changing agent as well as environment which acts as a positive feedback that further enhances diversity (Erwin, 2008). Indeed, such coevolutionary dynamics have also been a critical component to open-ended novelty search algorithms (Brant & Stanley, 2017; 2020; Wang et al., 2019). In these setups, the environment is defined defined by its own “genome” along with the agent itself, highlighting the extent to which the coevolutionary setting departs from a typical RL environment defined by an MDP. In such a coevolutionary setting where the state, action and reward landscapes (and therefore the solution landscape) are ever-shifting, the metaphor of “search”, with a putative end-point, proves less useful than “open-ended adaptation”.
@@ 2.2 Fitness as a Scalar but Multi-Objective Criterion
In RL problems too, there are many objectives that can be identified that aide in the achievement of the ultimate goal. The utility of each of these objectives may vary in degree across time, context, or even depend on history, but multiple objectives can also be simultaneously relevant at any given point in time. The main claim of the scalar reward hypothesis is that it is possible to distil all relevant objectives necessary for a given task, into a single scalar reward value (Silver et al., 2021). Even if the scalar reward hypothesis held true, and there were no proven bounds to its universality (Bowling et al., 2023), the question remains as to where an RL agent gets its reward objectives from. In artificial RL, the source of these objectives are ultimately rooted in the design decisions of the engineers that provide the auxiliary losses and tasks along with the “primary objective” of the task (Jaderberg et al., 2016). Some of these inductive biases are borrowed from psychology and biology of intrinsic motivation (IM) de Abril & Kanai (2018); Pathak et al. (2017); Houthooft et al. (2016); Kim et al. (2018); Pong et al. (2019); Leibfried & Grau-Moya (2019), while others may be generated adversarially (Campero et al., 2020) or even randomly without any a priori interpretable principle in mind (Osband et al., 2016; Eysenbach et al., 2018; Burda et al., 2018). In biological RL, on the other hand, the source of both intrinsic and extrinsically motivated behavior and their corresponding reward functions is ultimately attributed to cross-generational evolutionary mechanisms (Singh et al., 2010). In section 1.2, we demonstrated plausible mechanisms by which such evolutionary processes can take place within a single generation. This offers a plausible meta-learning mechanism by which multiple reward functions can be dynamically provisioned for the RL algorithms during the lifetime of the organism. Indeed, Czégel et al. (2021) speculate on the striatal reward system as a candidate location for assigning a scalar fitness value in addition to the state-action values of RL.
@@ 2.3 Evolution Does Not Explain Agency
It is tempting to attribute agency to evolution, both because evolution appears to be directed, and that it seems to provide a satisfying explanation as to where the reward functions of an organism come from Singh et al. (2010). However, the non-teleological nature of evolution is a subtle and largely uncontested characteristic of evolution Dawkins (1986); Gould (2002). More importantly, evolution as a theory, does not have any explanatory purchase in defining the replicative and homeostatic dynamics that facilitate it in the first place. Indeed, the roots of this argument go back to the days of the discoverers of evolutionary theory: Darwin and Wallace. While Darwin expressed doubts that the science of the day was able to conceive of such a theory for the origins of life (Darwin, 1863), Wallace resorted to more homuncular arguments in support of a “force vitale“ (Wallace, 1912).
@@ 3.1 Implications for Explore-Exploit
Exploration and Exploitation as Multiple Objectives The multi-objective nature of the explore-exploit problem is most clearly highlighted by Norman & Clune (2023) who point out that exploration and exploitation as objectives are fundamentally at odds with one another and that neither standard RL, meta-RL nor intrinsic motivation (IM) approaches are able to address this conflict appropriately. These authors distinguish between “coincidental” exploration strategies that these methods use and argue for “sacrificial” strategies that forgo short-term episode reward for longer term information gain. While methods like IM enable partial “sacrificial” exploration, Norman & Clune (2023) note that these strategies come at the cost of complete disengagement with the reward signal (even if temporary), leading to a number of known problems such as the noisy TV problem Ladosz et al. (2022). The solution offered by Norman & Clune (2023) is not evolutionary in nature; however, as discussed above, evolutionary learning strategies are adept at facilitating pursuit of multiple objectives. This is particularly true in the case of open-ended evolutionary algorithms where the search for novelty or diversity is in effect an exploratory strategy that continuously takes place alongside the search for high-quality individuals.

### v2

@@ Abstract
Three core tenets of reinforcement learning (RL)–concerning the definition of agency, the objective of learning, and the scope of the reward hypothesis–have been highlighted as key targets for conceptual revision, with major implications for theory and application. We propose a framework, inspired by open-ended evolutionary theory, to reconsider these three "dogmas." We revisit each assumption and address related concerns raised alongside them. To make our arguments relevant to RL as a model of biological learning, we first establish that evolutionary dynamics can plausibly operate within living brains over an individual’s lifetime, and are not confined to cross-generational processes. We begin by revisiting the second dogma, drawing on evolutionary insights to enrich the "adaptation-rather-than-search" view of learning. We then address the third dogma regarding the limits of the reward hypothesis, using analogies from evolutionary fitness to illuminate the scalar reward vs. multi-objective debate. After discussing practical implications for exploration in RL, we turn to the first–and arguably most fundamental–issue: the absence of a formal account of agency. We argue that unlike the other two problems, the evolutionary paradigm alone cannot resolve the agency question, though it gestures in a productive direction. We advocate integrating ideas from origins-of-life theory, where the thermodynamics of sustenance and replication offer promising foundations for understanding agency and resource-constrained reinforcement learning in biological systems.
@@ 1.1 Motivation & Overview
Finally, in section 3, we take a more practical turn and highlight areas where this theoretical framing can provide new tools for a practicing RL engineer. We specifically focus on the implications of the evolutionary framework on the problems of exploration-exploitation and resource-constrained reinforcement learning.
@@ 1.2 Evolution in the Brain
One of the leading figures in establishing the clonal selection theory of learning by the immune system, Gerald Edelman, was also one of the early advocates of a similar mechanism of learning by the nervous system (Edelman, 1993). This research program came to be known as neural Darwinism, and despite the pursuit of complementary ideas in a similar vein (Calvin, 1998), its adoption has been limited. While the volume of research on experience-guided learning via neuro-plasticity has catapulted over the years, research on the neural mechanisms and correlates of such neural Darwinism remain sparse.
@@ 2.1 Adaptation Versus Search
We will revisit the second point in section 2.2 when we discuss fitness in relation to scalar and multi-objective optimization. It is however constructive to elaborate on the first point and explain the mechanism with which novelty and diversity are achieved in these algorithms. Instead of “selection pressure”, the biological inspiration for novelty search is the notion of “speciation” or “niching” (Stanley & Miikkulainen, 2002; Mahfoud, 1996). This concept captures the tendency of organisms to found a new niche and thereby “avoid competition and exploit untapped resources” (Lehman & Stanley, 2010). QD algorithms encourage a diverse “archive” of solutions, by limiting selection pressure to individuals within the same niche, and avoiding premature selection against candidates that inhabit novel behavioral or phenotypic characteristics. An adjacent concept called “niche construction” (Odling-Smee et al., 2003; 2013) further enriches the adaptational view of learning in contrast to the search metaphor: organisms not only occupy a niche, but actively modify them. This results in coevolutionary dynamics between a changing agent as well as environment which acts as a positive feedback that further enhances diversity (Erwin, 2008). Indeed, such coevolutionary dynamics have also been a critical component to open-ended novelty search algorithms (Brant & Stanley, 2017; 2020; Wang et al., 2019). In these setups, the environment is defined by its own “genome” along with the agent itself, highlighting the extent to which the coevolutionary setting departs from a typical RL environment defined by an MDP. In such a coevolutionary setting where the state, action and reward landscapes (and therefore the solution landscape) are ever-shifting, the metaphor of “search”, with a putative end-point, proves less useful than “open-ended adaptation”.
@@ 2.2 Fitness as a Scalar but Multi-Objective Criterion
In RL problems too, there are many objectives that can be identified that aide in the achievement of the ultimate goal. The utility of each of these objectives may vary in degree across time, context, or even depend on history, but multiple objectives can also be simultaneously relevant at any given point in time. The main claim of the scalar reward hypothesis is that it is possible to distill all relevant objectives necessary for a given task, into a single scalar reward value (Silver et al., 2021). Even if the scalar reward hypothesis held true, and there were no proven bounds to its universality (Bowling et al., 2023), the question remains as to where an RL agent gets its reward objectives from. In artificial RL, the source of these objectives are ultimately rooted in the design decisions of the engineers that provide the auxiliary losses and tasks along with the “primary objective” of the task (Jaderberg et al., 2016). Some of these inductive biases are borrowed from psychology and biology of intrinsic motivation (IM) de Abril & Kanai (2018); Pathak et al. (2017); Houthooft et al. (2016); Kim et al. (2018); Pong et al. (2019); Leibfried & Grau-Moya (2019), while others may be generated adversarially (Campero et al., 2020) or even randomly without any a priori interpretable principle in mind (Osband et al., 2016; Eysenbach et al., 2018; Burda et al., 2018). In biological RL, on the other hand, the source of both intrinsic and extrinsically motivated behavior and their corresponding reward functions is ultimately attributed to cross-generational evolutionary mechanisms (Singh et al., 2010). In section 1.2, we demonstrated plausible mechanisms by which such evolutionary processes can take place within a single generation. This offers a plausible meta-learning mechanism by which multiple reward functions can be dynamically provisioned for the RL algorithms during the lifetime of the organism. Indeed, Czégel et al. (2021) speculate on the striatal reward system as a candidate location for assigning a scalar fitness value in addition to the state-action values of RL.
@@ 2.3 Evolution Does Not Explain Agency
It is tempting to attribute agency to evolution, both because evolution appears to be directed, and that it seems to provide a satisfying explanation as to where the reward functions of an organism come from (Singh et al., 2010). However, the non-teleological nature of evolution is a subtle and largely uncontested characteristic of evolution Dawkins (1986); Gould (2002). More importantly, evolution as a theory, does not have any explanatory purchase in defining the replicative and homeostatic dynamics that facilitate it in the first place. Indeed, the roots of this argument go back to the days of the discoverers of evolutionary theory: Darwin and Wallace. While Darwin expressed doubts that the science of the day was able to conceive of such a theory for the origins of life (Darwin, 1863), Wallace resorted to more homuncular arguments in support of a “force vitale“ (Wallace, 1912).
@@ 3.1 Implications for Explore-Exploit
Exploration and Exploitation as Multiple Objectives The multi-objective nature of the explore-exploit problem is most clearly highlighted by Norman & Clune (2023) who point out that exploration and exploitation as objectives are fundamentally at odds with one another and that neither standard RL, meta-RL nor intrinsic motivation (IM) approaches are able to address this conflict appropriately. These authors distinguish between “coincidental” exploration strategies that these methods use and argue for “sacrificial” strategies that forgo short-term episode reward for longer term information gain. While methods like IM enable partial “sacrificial” exploration, Norman & Clune (2023) note that these strategies come at the cost of complete disengagement with the reward signal (even if temporary), leading to a number of known problems such as the noisy TV problem (Ladosz et al., 2022). The solution offered by Norman & Clune (2023) is not evolutionary in nature; however, as discussed above, evolutionary learning strategies are adept at facilitating pursuit of multiple objectives. This is particularly true in the case of open-ended evolutionary algorithms where the search for novelty or diversity is in effect an exploratory strategy that continuously takes place alongside the search for high-quality individuals.